# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ B.2.5 Proof of Proposition 7
$\frac{\widehat{\theta}(n-\widehat{\theta})}{n}\leq\frac{\Delta}{4},\text{ \ i.e. \ }\frac{n}{\widehat{\theta}(n-\widehat{\theta})}\geq\frac{4}{\Delta}.$
@@
$\lesssim$ $\;\log^{2}(n)\cdot C_{1}\exp(-3\log\log(n))\lesssim C_{1}\log^{-1}(n),$
provided that $C\geq 12C_{0}$. Since $\mathsf{P}(\mathcal{A}_{1}\cap\mathcal{A}_{2})\geq 1-c_{2}\log^{-1}(n)$ with some $c_{2}$ that depends on $C_{1}$, we have
$\mathsf{P}\left[U_{1}\geq\frac{C\|\bm{\Sigma}\|\Psi^{2}\sqrt{p\log\log(n)}}{\Delta}\right]\lesssim\log^{-1}(n)$
@@ B.2.6 Supporting lemmas
Lemma B.5 (Lemma B.5 of Cho and Owens,, 2024).
@@
Lemma B.7 (Lemma B.1 of Verzelen and Gassiat,, 2018).
@@
See also Koltchinskii and Lounici, (2017) for a systematic treatment of this type of concentration inequalities.
Lemma B.8 (Lemma C.1 of Verzelen and Gassiat,, 2018).
@@
Lemma B.9 (Equation (27) of Verzelen and Gassiat,, 2018).
@@
Lemma B.10 (Lemma B.2 of Verzelen and Gassiat,, 2018).
@@
While the lemma is stated slightly more generally than Lemma B.2 of Verzelen and Gassiat, (2018), the proof takes analogous arguments and thus is omitted; see also the proof of Lemma B.11. ∎
@@
[…]
@@
The proof largely follows the proof of Lemma B.2 of Verzelen and Gassiat, (2018), which in turn relies on the deviation inequality derived for Gaussian chaos in Theorem 1.3 of Adamczak and Wolff, (2015).
@@
[…]
where $\mathbf{z}^{(l)}$ is a $|J_{l}|$-indexed matrix. Then, $f(\mathbf{Z})$ is a polynomial of degree $D=4$ of the $q=np$ independent standard Gaussian random variables $Z_{ti}$’s, to which Theorem 1.3 of Adamczak and Wolff, (2015) is applicable, such that with some constant $C_{0}>0$,
@@
where $\partial^{d}f$ denotes the $d$-th derivative of $f$; we later set
@@
Denoting by $\mathbf{C}=[c_{(t_{1},i_{1}),(t_{2},i_{2}),(t_{3},i_{3}),(t_{4},i_{4})}]$ with $(t_{l},i_{l})\in[n]\times[p]$, we derive a bound on $\|\mathsf{E}(\partial^{4}f(\mathbf{Z}))\|_{\mathcal{J}}$ from that on $\|\mathbf{C}\|_{\mathcal{J}}$ by triangle inequality, as $[\mathsf{E}(\partial^{4}f(\mathbf{Z}))]_{\mathbf{i}_{1},\mathbf{i}_{2},\mathbf{i}_{3},\mathbf{i}_{4}}=\sum_{\pi}C_{\mathbf{i}_{\pi(1)},\mathbf{i}_{\pi(2)},\mathbf{i}_{\pi(3)},\mathbf{i}_{\pi(4)}}$, where the sum is taken over all permutations $\pi$ of [4][4][ 4 ]. When $\mathcal{J}=\{[4]\}$ (hence $|\mathcal{J}|=1$), it holds
@@
$c^{\prime}_{(t_{1},i_{1}),(t_{2},i_{2})}=\left\{\begin{array}[]{ll}(n-k)(a_{i_{1}}b_{i_{2}}\mu_{i_{2}}+a_{i_{2}}b_{i_{1}}\mu_{i_{1}})&\text{when \ }1\leq t_{1}=t_{2}\leq k,\\ k(a_{i_{1}}b_{i_{2}}\mu_{i_{1}}+a_{i_{2}}b_{i_{1}}\mu_{i_{2}})&\text{when \ }k+1\leq t_{1}=t_{2}\leq n,\\ 0&\text{when \ }t_{1}\neq t_{2}.\\ \end{array}\right.$
@@
B.3 Proofs of the results in Section 4
@@ B.3.1 Proof of Proposition 8
$\bar{\mathcal{E}}$ $:=\bigcap_{k\in\mathcal{D}\cap\bar{\mathcal{Q}}}\left\{|\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)-\bar{\mathbf{f}}_{k}|_{\infty}\leq\bar{C}_{0}^{\prime}\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\log(p\log(n))}\right\},$ (B.19a)
@@
where $\mathcal{D}$ denotes the dyadic grid in Algorithm 1 and $\bar{\mathcal{Q}}$ and $\mathcal{Q}$ deterministic grids that depend only on $\bar{\mathbf{f}}_{k}$ or $f_{k}$, respectively (see Step 2 in the proof of Theorem 5), and $\bar{C}_{0}^{\prime}$ and $C_{0}^{\prime}$ some universal constants that depend on $\bar{C}_{0}$ and $C_{0}$, respectively. By Propositions B.4 and 4, and from that $|\mathcal{D}|\leq 2\log_{2}(n)$ and $|\bar{\mathcal{Q}}|,|\mathcal{Q}|\leq 16\log^{2}(n)$, it follows that $\mathsf{P}(\bar{\mathcal{E}}\cap\mathcal{E})\geq 1-\bar{c}_{2}(p\log(n))^{-1}-c_{2}\log^{-1}(n)$ with $\bar{c}_{2}$ and $c_{2}$ depending on $\bar{C}_{1}$ and $C_{1}$, respectively. Below we condition our arguments on $\bar{\mathcal{E}}\cap\mathcal{E}$.
@@
$\frac{|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}\Delta}{\log(p\log(n))}\lesssim\|\bm{\Sigma}\|\Psi^{2}\lesssim\frac{|\bm{\Sigma}\bm{\delta}|_{2}^{2}\Delta}{\sqrt{p\log\log(n)}},$
@@
$\zeta_{\text{\tiny M}}<\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}\leq|\bm{\Sigma}\bm{\delta}|_{\infty}\sqrt{\Delta}+\bar{C}_{0}^{\prime}\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\log(p\log(n))},$ (B.20)
@@
$\sqrt{\frac{1-\eta}{2(1+\eta)}}|\bm{\Sigma}\bm{\delta}|_{\infty}\sqrt{\Delta}-\bar{C}_{0}^{\prime}\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\log(p\log(n))}\leq\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}.$ (B.21)
@@
B.3.2 Proof of Proposition 9
@@
$\left|\widehat{\bm{\delta}}-\bm{\delta}(\widehat{\theta}_{\text{\tiny O}})\right|_{1}\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\mathfrak{s}_{\delta}\sqrt{\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}},$
@@
$\left|\bm{\delta}(\widehat{\theta}_{\text{\tiny O}})-\bm{\delta}\right|_{1}\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi}{\underline{\sigma}\sqrt{\Delta}}\min\left\{\mathfrak{s}_{\delta}\sqrt{\log(p\log(n))},\,\sqrt{\mathfrak{s}_{\delta}}\bigl{(}p\log\log(n)\bigr{)}^{1/4}\right\},$
which leads to
@@
$\lesssim$ $\,\frac{\mathfrak{s}_{\delta}\|\bm{\Sigma}\|\Psi^{2}\log(p\vee n)}{\underline{\sigma}\sqrt{\Delta}}.$
@@
$\lesssim$ $\,|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}\Psi\sqrt{\log(p\vee n)}+\frac{\mathfrak{s}_{\delta}\|\bm{\Sigma}\|\Psi^{2}\log(p\vee n)}{\underline{\sigma}\sqrt{\Delta}}.$ (B.25)
@@
$\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}-(\widehat{\bm{\delta}}-\bm{\delta})^{\top}\bm{\Sigma}\bm{\delta}$ $\geq\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}-\|\bm{\Sigma}^{1/2}\||\bm{\Sigma}^{1/2}\bm{\delta}|_{2}|\widehat{\bm{\delta}}-\bm{\delta}|_{2}$
@@
as long as $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta\geq 4C\underline{\sigma}^{-2}\|\bm{\Sigma}\|^{2}\Psi^{2}\mathfrak{s}_{\delta}\log(p\vee n)$, a condition readily met under (12) for large enough $\widetilde{c}_{0}$. Given these observations, we first show that $|\widehat{\theta}_{\text{\tiny R}}-\theta|\leq\Delta/2$. This follows from the fact that for any $\varpi_{\text{\tiny R}}<k<\theta-\Delta/2$, we have from (B.25) and (B.26),
@@
==:= : $\,U_{1}+U_{2}-U_{3}.$
Using the arguments leading to (B.25) in combination with those adopted in the proof of Theorem 2 of Cho et al., (2025), we have
@@
while from that $|\widehat{\theta}_{\text{\tiny R}}-\theta|\leq\Delta/2$, (B.26) and Lemma 7 of Wang and Samworth, (2018),
@@
$|\displaystyle|\widehat{\theta}_{\text{\tiny R}}-\theta|$$\lesssim|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{-2}\Psi^{2}\log(p\vee n)\max\left\{1,\,\frac{\|\bm{\Sigma}\|^{2}\Psi^{2}\mathfrak{s}_{\delta}^{2}\log(p\vee n)}{\underline{\sigma}^{2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta}\right\}$

### v3

@@ B.2.5 Proof of Proposition 7
[…]
@@
$\lesssim$ $\;\log^{2}(n)\cdot C_{1}\exp(-3\log\log(n))\lesssim C_{1}\bigl(\log(n)\bigr)^{-1},$
provided that $C\geq 12C_{0}$. Since $\mathsf{P}(\mathcal{A}_{1}\cap\mathcal{A}_{2})\geq 1-c_{2}\bigl(\log(n)\bigr)^{-1}$ with some $c_{2}$ that depends on $C_{1}$, we have
$\mathsf{P}\left[U_{1}\geq\frac{C\|\bm{\Sigma}\|\Psi^{2}\sqrt{p\log\log(n)}}{\Delta}\right]\lesssim\bigl(\log(n)\bigr)^{-1}$
@@ B.2.6 Supporting lemmas
Lemma B.5 (Lemma B.5 of Cho and Owens, 2024).
@@
Lemma B.7 (Lemma B.1 of Verzelen and Gassiat, 2018).
@@
See also Koltchinskii and Lounici (2017) for a systematic treatment of this type of concentration inequalities.
Lemma B.8 (Lemma C.1 of Verzelen and Gassiat, 2018).
@@
Lemma B.9 (Equation (27) of Verzelen and Gassiat, 2018).
@@
Lemma B.10 (Lemma B.2 of Verzelen and Gassiat, 2018).
@@
While the lemma is stated slightly more generally than Lemma B.2 of Verzelen and Gassiat (2018), the proof takes analogous arguments and thus is omitted; see also the proof of Lemma B.11. ∎
@@
Let $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})$ for $t\in[n]$, $p\geq n$ and $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\top}$. Consider some $0<k<n$. For all $z\geq 0$, any $\bm{\alpha},\bm{\beta}\in\mathbb{R}^{p}$ and some constant $C>0$, we have
@@
The proof largely follows the proof of Lemma B.2 of Verzelen and Gassiat (2018), which in turn relies on the deviation inequality derived for Gaussian chaos in Theorem 1.3 of Adamczak and Wolff (2015).
@@
To proceed further, we need some notation. For positive integers $d$ and $q$, we denote a $d$-indexed matrix $\mathbf{C}=(c_{i_{1},\ldots,i_{d}})_{i_{1},\ldots,i_{d}=1}^{q}$; also we write $\mathbf{i}=(i_{1},\ldots,i_{d})\in[q]^{d}$ and $\mathbf{i}_{J}=(i_{l},\,l\in J)$ for $J\subseteq[d]$. Also, let $\mathcal{P}_{d}$ the set of partitions of $[d]$ into non-empty disjoint subsets. Given a partition $\mathcal{J}=\{J_{1},\ldots,J_{\ell}\}$, we define the tensor product norm:
$\|\mathbf{C}\|_{\mathcal{J}}:=\sup\left\{\sum_{\mathbf{i}\in[q]^{d}}c_{\mathbf{i}}\prod_{l=1}^{\ell}\mathbf{z}^{(l)}_{\mathbf{i}_{J_{l}}}:\,\|\mathbf{z}^{(l)}\|_{F}\leq 1,\,1\leq l\leq\ell\right\},$
where $\mathbf{z}^{(l)}$ is a $|J_{l}|$-indexed matrix. Then, $f(\mathbf{Z})$ is a polynomial of degree $D=4$ of the $q=np$ independent standard Gaussian random variables $Z_{ti}$’s, to which Theorem 1.3 of Adamczak and Wolff (2015) is applicable, such that with some constant $C_{0}>0$,
@@
where $\partial^{d}f$ denotes the $d$th derivative of $f$; we later set
@@
Denoting by $\mathbf{C}=[c_{(t_{1},i_{1}),(t_{2},i_{2}),(t_{3},i_{3}),(t_{4},i_{4})}]$ with $(t_{l},i_{l})\in[n]\times[p]$, we derive a bound on $\|\mathsf{E}(\partial^{4}f(\mathbf{Z}))\|_{\mathcal{J}}$ from that on $\|\mathbf{C}\|_{\mathcal{J}}$ by triangle inequality, as $\mathsf{E}(\partial^{4}f(\mathbf{Z}))]_{\mathbf{i}_{1},\mathbf{i}_{2},\mathbf{i}_{3},\mathbf{i}_{4}}=\sum_{\pi}C_{\mathbf{i}_{\pi(1)},\mathbf{i}_{\pi(2)},\mathbf{i}_{\pi(3)},\mathbf{i}_{\pi(4)}}$, where the sum is taken over all permutations $\pi$ of $[4]$. When $\mathcal{J}=\{[4]\}$ (hence $|\mathcal{J}|=1$), it holds
@@
[…]
@@
B.3 Proofs for the results in Section 4
@@ B.3.1 Proof of Proposition 8
$\bar{\mathcal{E}}$ $:=\bigcap_{k\in\mathcal{D}\cap\bar{\mathcal{Q}}}\left\{|\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)-\bar{\mathbf{f}}_{k}|_{\infty}\leq\bar{C}_{0}^{\prime}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\log(p\log(n))}\right\},$ (B.17a)
@@
where $\mathcal{D}$ denotes the dyadic grid in Algorithm 1 and $\bar{\mathcal{Q}}$ and $\mathcal{Q}$ deterministic grids that depend only on $\bar{\mathbf{f}}_{k}$ or $f_{k}$, respectively (see Step 2 in the proof of Theorem 5), and $\bar{C}_{0}^{\prime}$ and $C_{0}^{\prime}$ some universal constants that depend on $\bar{C}_{0}$ and $C_{0}$, respectively. By Propositions B.4 and 4, and from that $|\mathcal{D}|\leq 2\log_{2}(n)$ and $|\bar{\mathcal{Q}}|,|\mathcal{Q}|\leq 16\log^{2}(n)$, it follows that $\mathsf{P}(\bar{\mathcal{E}}\cap\mathcal{E})\geq 1-\bar{c}_{2}(p\log(n))^{-1}-c_{2}\bigl(\log(n)\bigr)^{-1}$ with $\bar{c}_{2}$ and $c_{2}$ depending on $\bar{C}_{1}$ and $C_{1}$, respectively. Below we condition our arguments on $\bar{\mathcal{E}}\cap\mathcal{E}$.
@@
$\frac{|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}\Delta}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}\log(p\log(n))}\lesssim\Psi^{2}\lesssim\frac{|\bm{\Sigma}\bm{\delta}|_{2}^{2}\Delta}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bm{\Sigma}\|}\sqrt{p\log\log(n)}},$
@@
$\zeta_{\text{\tiny M}}<\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}\leq|\bm{\Sigma}\bm{\delta}|_{\infty}\sqrt{\Delta}+\bar{C}_{0}^{\prime}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\log(p\log(n))},$ (B.18)
@@
$\sqrt{\frac{1-\eta}{2(1+\eta)}}|\bm{\Sigma}\bm{\delta}|_{\infty}\sqrt{\Delta}-\bar{C}_{0}^{\prime}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\log(p\log(n))}\leq\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}.$ (B.19)
@@
B.3.2 Proof of Corollary 9
We refer to Lemma 2 of Moen (2024) for the proof of the probabilistic statements below.
Denoting the diagonal entries of $\bm{\Sigma}$ by $\Sigma_{ii}$, there exists $C\in(0,\infty)$ such that
$\mathsf{P}\left(\max_{i\in[p]}\left\lvert\frac{\widehat{\Sigma}_{ii}}{\Sigma_{ii}}-1\right\rvert\leq C\sqrt{\frac{\log(p)}{n}}\right)\geq 1-p^{-1},$
$\mathsf{P}\left(\bigcap_{\ell\in\{1,2\}}\bigcap_{t\in\mathcal{T}}\left\{\left|\widehat{\Psi}^{2}_{\ell,t}-\mathsf{E}(\widehat{\Psi}^{2}_{\ell,t})\right|\leq\frac{\Psi^{2}}{2}\sqrt{\frac{\log\log(n)}{t}}\right\}\right)\geq 1-C\log^{-1}(n).$ (B.21)
Next, let us set $\widehat{\Psi}^{2}=\max_{t\in\mathcal{T}}(\widehat{\Psi}^{2}_{1,t}\vee\widehat{\Psi}^{2}_{2,t})$. We observe that
$\mathsf{E}(\widehat{\Psi}^{2}_{1,t})$ $=\frac{t\wedge\theta}{t}|\bm{\Sigma}\bm{\beta}_{0}|_{2}^{2}+\frac{(t-\theta)\vee 0}{t}|\bm{\Sigma}\bm{\beta}_{1}|_{2}^{2}+\sigma^{2}\leq 3\Psi^{2},\text{ \ and}$
$\mathsf{E}(\widehat{\Psi}^{2}_{2,t})$ $=\frac{t\wedge(n-\theta)}{t}|\bm{\Sigma}\bm{\beta}_{1}|_{2}^{2}+\frac{(t-n+\theta)\vee 0}{t}|\bm{\Sigma}\bm{\beta}_{0}|_{2}^{2}+\sigma^{2}\leq 3\Psi^{2},$
which leads to $\widehat{\Psi}^{2}\lesssim\Psi^{2}$ with probability tending to one thanks to (B.21) and the construction of $\mathcal{T}$. WLOG, let us suppose that $|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2}\geq|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2}$. Then, for any $t\leq\min(n/2,n-\theta)$, we have $\widehat{\Psi}^{2}_{2,t}\geq\Psi^{2}/2$ uniformly with probability tending to one. Summarising these observations, we have
$\mathsf{P}\left(M^{-1}\leq\frac{\widehat{\Psi}^{2}}{\Psi^{2}}\leq M\right)\geq 1-C\log^{-1}(n),$
where $M\in(1,\infty)$ is an absolute constant.
Finally, supposing that $p\leq c_{*}n$ for a constant $c_{*}\in(0,\infty)$, there exists $C\in(0,\infty)$ that depends only on $c_{*}$, which gives
$\mathsf{P}\left(\|\widehat{\bm{\Sigma}}-\bm{\Sigma}\|\leq C\|\bm{\Sigma}\|\sqrt{\frac{p}{n}}\right)\geq 1-e^{-p}.$
Summarising the above concludes the proof.
B.3.3 Proof of Proposition 10
@@
$\left|\widehat{\bm{\delta}}-\bm{\delta}(\widehat{\theta}_{\text{\tiny O}})\right|_{1}\lesssim\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\mathfrak{s}_{\delta}\sqrt{\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}},$
@@
$\left|\bm{\delta}(\widehat{\theta}_{\text{\tiny O}})-\bm{\delta}\right|_{1}\lesssim\frac{\Psi}{\underline{\sigma}\sqrt{\Delta}}\min\left\{\mathfrak{s}_{\delta}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\sqrt{\log(p\log(n))},\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bm{\Sigma}\|^{1/2}}\sqrt{\mathfrak{s}_{\delta}}\bigl(p\log\log(n)\bigr)^{1/4}\right\}$
which, combined with that $\|\bm{\Sigma}\|^{1/2}\geq\sigma_{X}$, leads to
@@
$\lesssim$ $\,\frac{\mathfrak{s}_{\delta}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bm{\Sigma}\|^{1/2}\sigma_{X}}\Psi^{2}\log(p\vee n)}{\underline{\sigma}\sqrt{\Delta}}.$
@@
$\lesssim$ $\,|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}\Psi\sqrt{\log(p\vee n)}+\frac{\mathfrak{s}_{\delta}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bm{\Sigma}\|^{1/2}\sigma_{X}}\Psi^{2}\log(p\vee n)}{\underline{\sigma}\sqrt{\Delta}}.$ (B.24)
@@
$\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}-(\widehat{\bm{\delta}}-\bm{\delta})^{\top}\bm{\Sigma}\bm{\delta}$ $\geq\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}-\|\bm{\Sigma}\|^{1/2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}|\widehat{\bm{\delta}}-\bm{\delta}|_{2}$
@@
as long as $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta\geq 4C^{2}\underline{\sigma}^{-2}\|\bm{\Sigma}\|^{2}\Psi^{2}\mathfrak{s}_{\delta}\log(p\vee n)$, a condition readily met under (12) for large enough $\widetilde{c}_{0}$. Given these observations, we first show that $|\widehat{\theta}_{\text{\tiny R}}-\theta|\leq\Delta/2$. This follows from the fact that for any $\varpi_{\text{\tiny R}}<k<\theta-\Delta/2$, we have from (B.24) and (B.25),
@@
$=:$ $\,U_{1}+U_{2}-U_{3}.$
Using the arguments leading to (B.24) in combination with those adopted in the proof of Theorem 2 of Cho et al. (2025), we have
@@
while from that $|\widehat{\theta}_{\text{\tiny R}}-\theta|\leq\Delta/2$, (B.25) and Lemma 7 of Wang and Samworth (2018),
@@
$|\widehat{\theta}_{\text{\tiny R}}-\theta|$ $\lesssim|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{-2}\Psi^{2}\log(p\vee n)\max\left\{1,\,\frac{\|\bm{\Sigma}\|^{2}\Psi^{2}\mathfrak{s}_{\delta}^{2}\log(p\vee n)}{\underline{\sigma}^{2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta}\right\}$
@@
B.3.4 Minimax localisation rate under strong signal strength
The localisation rate obtained in Proposition 10 is minimax optimal up to a logarithmic factor. This can be seen by projecting the model onto the oracle direction $\bm{\delta}$ and considering the resulting univariate mean change problem for $\{\bm{\delta}^{\top}\mathbf{x}_{t}Y_{t}\}_{t\in[n]}$, for which the optimal localisation rate for this problem is well understood, see e.g. Verzelen et al. (2023). Regarding the problem under (1) as that of detecting a change point in the mean of the $p$-variate sequence $\{\mathbf{x}_{t}Y_{t}\}_{t\in[n]}$, such an oracle approach can be shown to attain the maximum high-dimensional efficiency as defined by Aston and Kirch (2018).
We now provide a formal argument using standard lower bound techniques based on two hypotheses. Consider the single change point model under (1), and define the model space
$\mathcal{P}_{\bm{\Sigma},\sigma}^{\mathfrak{s}_{\delta},n,p}(\tau)=\biggl\{\mathsf{P}_{\theta,\bm{\beta}_{0},\bm{\beta}_{1}}\;:\;\{(Y_{t},\mathbf{x}_{t})\}_{t\in[n]}\sim\mathsf{P}_{\theta,\bm{\beta}_{0},\bm{\beta}_{1}}\text{ \ such that \ }\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})\text{ and}\\ \text{independently \ }\varepsilon_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2}),\;Y_{t}=\mathbf{x}_{t}^{\top}\bigl(\bm{\beta}_{0}\cdot\mathbb{I}_{\{t\leq\theta\}}+\bm{\beta}_{1}\cdot\mathbb{I}_{\{t>\theta\}}\bigr)+\varepsilon_{t},\\ \bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0},\,\Delta=\min(\theta,n-\theta),\,\,\Psi=\max(|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2},|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2},\sigma),\\ \lvert\bm{\delta}\rvert_{0}=\mathfrak{s}_{\delta}\text{ \ and \ }\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}\cdot\Delta\geq\tau\Psi^{2}\mathfrak{s}_{\delta}^{2}\log(p\vee n)\biggr\}$
for some constant $\tau>0$, with $\bm{\Sigma}$ positive definite and $\sigma>0$.
Lemma B.13.
Let $\tau$ be either a sufficiently large constant (in particular, $\tau\geq\widetilde{c}_{0}\underline{\sigma}^{-2}\|\bm{\Sigma}\|^{2}$) or satisfy $\tau\to\infty$ (at an arbitrary rate) as $n\to\infty$. If
$n\geq 16\tau\mathfrak{s}_{\delta}^{2}\log(p\vee n),$ (B.26)
then
$\inf_{\widehat{\theta}}\sup_{\mathsf{P}\in\mathcal{P}_{\bm{\Sigma},\sigma}^{\mathfrak{s}_{\delta},n,p}(\tau)}\mathsf{P}\!\left(\left|\widehat{\theta}\bigl((Y_{t},\mathbf{x}_{t})_{t\in[n]}\bigr)-\theta\right|\geq\frac{\Psi^{2}}{8\,\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}}\right)\geq\frac{1}{4},$
where the infimum is taken over all measurable functions $\widehat{\theta}:\mathbb{R}^{n\times(p+1)}\to[n]$.
Proof.
Consider the two hypotheses
$\mathsf{P}^{(0)}=\mathsf{P}_{n/4,\bm{\delta},\bm{0}}\quad\text{and}\quad\mathsf{P}^{(1)}=\mathsf{P}_{n/4+2\gamma,\bm{\delta},\bm{0}},$
where $\bm{\delta}$ satisfies $|\bm{\delta}|_{0}=\mathfrak{s}_{\delta}$ and $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}=\sigma^{2}/4$, with $\gamma\leq n/4$. In this setting, $\Psi^{2}=\sigma^{2}$. Condition (B.26) ensures that $\mathsf{P}^{(0)},\mathsf{P}^{(1)}\in\mathcal{P}_{\bm{\Sigma},\sigma}^{\mathfrak{s}_{\delta},n,p}(\tau)$, and therefore
$\inf_{\widehat{\theta}}\;\sup_{\mathsf{P}\in\mathcal{P}_{\bm{\Sigma},\sigma}^{\mathfrak{s}_{\delta},n,p}(\tau)}\mathsf{P}\!\left(\bigl|\widehat{\theta}-\theta(\mathsf{P})\bigr|\geq\gamma\right)\geq\frac{1-\sqrt{\alpha/2}}{2},$
where $\alpha=\chi^{2}(\mathsf{P}^{(1)},\mathsf{P}^{(0)})$. By Lemma B.2,
$\alpha+1=\left(\mathsf{E}_{\mathsf{P}_{\bm{0}}}\!\left[\left(\frac{\mathrm{d}\mathsf{P}_{\bm{\delta}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)^{2}\right]\right)^{2\gamma}=2^{\gamma}.$
Choosing $\gamma:={\Psi^{2}}/({8\,\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}})=1/2$ yields $\alpha=\sqrt{2}-1\leq 1/2$, and hence
$\frac{1-\sqrt{\alpha/2}}{2}\geq\frac{1}{4},$
which completes the proof. ∎